OpenStudy (anonymous):

Object A is moving down a frictionless path at a velocity of 5 m/s. If it hits and sticks to object B, which has twice the mass of object A, what would the resulting velocity be?

OpenStudy (anonymous):

Using conservation of momentum, \[m _{1}v _{1}=m _{2}v _{2}\] Since we know that \[m _{2}=2m\] we can rewrite the above equation as \[m _{1}*5m/s=2m _{1}v _{2}\] Finally, we can divide 2*m1 to isolate v2. So,\[v_{2} = (5/2) m/s\]

OpenStudy (anonymous):

Thanks!